OpenStudy (anonymous):

If a ball is thrown directly upward with a velocity of 20 ft/s, its height (in feet) after t seconds is given by y = 20t − 16t2. What is the maximum height attained by the ball? (Round your answer to the nearest whole number.)

OpenStudy (anonymous):

y = 20t − 16t^2

OpenStudy (anonymous):

max is at the second coordinate of the vertex the first coordinate of the vertex is \(t=-\frac{b}{2a}\) which in your case is \(-\frac{20}{2\times (-16)}=\frac{5}{8}\)

OpenStudy (anonymous):

replace \(t\) by \(\frac{5}{8}\) to get the maximum height use a calculator

OpenStudy (anonymous):

i get this http://www.wolframalpha.com/input/?i=20*%285%2F8%29-16%285%2F8%29^2

OpenStudy (anonymous):

I just got that. thanks :)